In a recent year the total sales at Restaurant A were $250,000 more than Restaurant B. Total sales at the two stores were $625,000,000. What were the sales for the year at each store?
Answer by checkley77(12844)

X+(X+250,000)=625,000,000
2X=625,000,000-250,000
2X=624,750,000
X=624,750,000/2
X=312,375,000 ANS. FOR RESTAURANT B.
312,375,000+250,000=312,625,000 ANS. FOR RESTAURANT A.